Grade 12th passAlgebra

Convert each part of the ratio to percentage:
3 : 1

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Harshit Singh
5 Years ago
Dear Student

We have to find total parts by adding the given ratio = 3 + 1 = 4
1stpart = 3⁄4 = (3⁄4) × 100 %
= 3 × 25%
= 75%
2ndpart = 1⁄4 = (1⁄4) × 100%
= 1 × 25 = 25%

Thanks